OpenStudy (puppylover100):

Factor completely 50a2b5 - 35a4b3 + 5a3b4 A. 5a2b3(10b2 - 7a2 + ab) B. 10b2 - 7a2 + ab C. a2b3(50b2 - 35a2 + 5ab) D. 5(10a2b5 - 7a4b3 + a3b4)

OpenStudy (puppylover100):

Will medal

jagr2713 (jagr2713):

Hey what do you think we do first

jagr2713 (jagr2713):

What number can go into both 50,35,5?

jagr2713 (jagr2713):

you there @Puppylover100

OpenStudy (puppylover100):

5

jagr2713 (jagr2713):

Ok 5 do you think we can factor anything else out like variables ?

OpenStudy (puppylover100):

Well there is a^2 b^5 a^4 b^3 a^3 b^4 the two most common numbers are 3 and 4 2 can go into four ut not 5 or 3

jagr2713 (jagr2713):

so from looking at the answer choices, what do you think th answer is?

OpenStudy (puppylover100):

D

jagr2713 (jagr2713):

D??? do you know how to factor?

OpenStudy (puppylover100):

Why wouldn't be D? If you muliply everything by 5 you would get the original problem

jagr2713 (jagr2713):

No we have to divide

jagr2713 (jagr2713):

ok 5 goes into both correct? So we divide 5 by the 50,35,5

OpenStudy (puppylover100):

5 goes into 50 ten times, 35 seven times, and 5 once

jagr2713 (jagr2713):



jagr2713 (jagr2713):

a and b goes into all 3 terms so we find the smallest of both A and B so we have a2 and b3 and we subtract it from all 3 terms

OpenStudy (puppylover100):

ok

jagr2713 (jagr2713):

so 50/5=10 so we subtract 2-2=0 5-3=2

jagr2713 (jagr2713):

so what do we get

OpenStudy (puppylover100):

10b^2

jagr2713 (jagr2713):

correct because 2-2=0 so a cancels out :D how about the next term

OpenStudy (puppylover100):

7a^2

jagr2713 (jagr2713):

how about the next

OpenStudy (puppylover100):

ab

jagr2713 (jagr2713):

So the answer is?

OpenStudy (puppylover100):

10b^2-7a^2+ab

jagr2713 (jagr2713):

you forgot the something in the beginning

OpenStudy (puppylover100):

5(10b^2-7a^2+ab)

jagr2713 (jagr2713):

you forgot the variables

OpenStudy (puppylover100):

5a^2b^3(10b^2-7a^2+ab)

jagr2713 (jagr2713):

Correct :D

OpenStudy (puppylover100):

Thank you SO much

jagr2713 (jagr2713):

Anytime

jagr2713 (jagr2713):

\[ \huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\] if you want to know more about this site, Here is a tutorial about this site : http://openstudy.com/study#/updates/5525c0fce4b030ceb3f357e8

OpenStudy (puppylover100):

:)

OpenStudy (puppylover100):

Thanks

jagr2713 (jagr2713):

:D

OpenStudy (puppylover100):

:D